Tuesday, August 2, 2016

some observations about buoy status

The following is a brief status update based on a close examination of the data record, in a similar manner to an earlier update last May.
  • The analog RH/AirT sensor remains nonfunctional. Based on my direct observation of similar problems with the Calabash Caye buoy I believe this is probably a buoy wiring problem and not a failure of the instrument itself.
  • The EXO's CT2 sensor has performed well w.r.t. sea temperatures, although there continues to be reason to doubt the sensor's overlow salinity reports, even following a March recalibration. Also the EXO does not appear to have been calibrated again since March, unlike the BUTO1 EXO which was recalibrated in May when its BICs were installed.
  • Diagnostic data continue to suggest that the interior of the 'Main' canister is unacceptably moist (although this is even more seriously true of the BUTO1 buoy).
 (posted by Mike Jankulak)

Thursday, July 7, 2016

quick buoy update from Jonathan Gomez

[The following is another (slightly edited) email received from Jonathan Gomez on July 7, 2016. It contains information about both Tobago buoys, this one and the one at Buccoo Reef.]
Just a quick update on the CREWS. There are a couple issues we would like some assistance with.

1. Still no Dissolved oxygen or pH readings are showing up in the data stream from the EXO2 when connected to the CREWS. However, all sensors are showing up and giving readings when connected via Bluetooth when we are calibrating. Please note that the DO sensor was installed in error on the Speyside CREWS last December, but this has since been corrected.

2. We were unable to establish a connection to the CREWS using Loggernet when connected with the supplied umbilical cable. Thus, we were not able to install the software to add the new BIC instruments to the data stream.
(posted by Mike Jankulak)

Thursday, May 26, 2016

met data from this buoy are now feeding NDBC and the world

Last week I began feeding the 10-minute meteorological data from the two Tobago buoys to the National Data Buoy Center (NDBC).  A few days later NDBC notified me that those data feeds have been cleared for release to the Global Telecommunications Systems (GTS), which makes those data available for use by national weather services all over the world.

The Angel's Reef (Speyside) buoy, which we informally refer to as ARTO1, has officially been designated 42088, which is in keeping with the custom of giving all-numeric designations to buoys.  The NDBC home page for the Angel's Reef buoy may be found here:

http://www.ndbc.noaa.gov/station_page.php?station=42088

This link has also been added to the 'Links' section on the right-hand side of this blog.


regards,
Mike J+

Monday, May 23, 2016

news and photo from Jonathan Gomez

[Note: the following is a slightly edited version of an email received from Jonathan Gomez (who, along with Addison Titus, have been maintaining the two CREWS buoys in Tobago since Jahson Alemu's departure last March) on May 23, 2016. This post is back-dated to that date.]

Click on image to see original size.
I'll give you some updates on the two CREWS.  Myself and Addison removed the EXOs in March for servicing and calibration. Also, both CREWS's support lines and buoy were cleaned of fouling and inspected for wear. Also, in early May the BIC hardware was installed and we took the opportunity to service and calibrate the Buccoo EXO again. The software still needs to be installed for the BIC sensor to be added to the data stream however. We plan to do this in the next few weeks with guidance from Jon Fajans.

There are a couple issues with data from the EXOs however. For the Buccoo EXO, all parameters are functioning fine while doing the calibrations, but when connected to the CREWS we aren't seeing Dissolved Oxygen and pH in the data streams.

It also turned out there was a mix-up with the DO sensor. It was accidentally installed on the Speyside EXO during the December 2015 servicing. This with be corrected on our next trip mentioned above.

I think those are the main issues we have at the moment, if we have any others I'll let you know.

Both ADCPs are deployed and running. We downloaded the last set of data in March. I believe Jahson would do this every 6 months so we are continuing with that schedule for now.

I've attached two photos of the CREWS for you, 'Picture 303' is the Speyside one, and 'Picture 255' is the Buccoo one.
posted by Mike Jankulak

Thursday, May 19, 2016

status, based on examination of data stream

Since there has been little recent activity on this maintenance log, I've decided to post a few remarks based solely on my examination of the data stream from this station.

As a reminder, this station's recent history includes its annual equipment swapout (December 7th - 8th, 2015), at which time the EXO Sonde was reinstalled after been absent for warranty repairs since July 21st, 2015.  These repairs included an upgrade to a titanium bulkhead and the replacement of the CT module (conductivity/temperature sensor, which reports salinity as well) with an improved CT2 module.  It was hoped that the reengineered CT2 module would eliminate the many problems with failures of the temperature and salinity parameters experienced over this station's lifetime since late 2013.  After the 12/8/2015 swapout was completed, no further updates were received apart from an announcement on February 5th, 2016, that Jahson Alemu would no longer be our primary local contact for the Tobago buoys.

As of May 16th, 2016, the following observations can be made about this buoy's data stream:
  • The EXO Sonde was indeed reconnected during the annual swapout and started producing data on December 8th, 2015, after its 4.5 month absence.
  • The EXO Sonde went offline from March 18th to 23rd, 2016. My guess would be that the instrument was removed for five days during some kind of maintenance operation but it is possible that it simply failed.
  • Since last December, apart from the 5-day interruption in March, the EXO seems have produced consistent and reliable sea temperature data.
  • However the EXO's salinity data appears to be somewhat questionable. At first it reported in the neighborhood of 36 PSU, which then rapidly and steadily fell by more than 2 PSU over only 7 days from January 26th to February 2nd, 2016.  After the 5-day outage in March its reports remained low, maxing out at about 33 PSU but falling in recent days below 31 PSU. As of this writing I am not confident that the conductivity (salinity) part of this sensor is functioning properly. This may indicate that the CT2 module has not corrected all of the problems of the CT module, or it may suggest that the required monthly on-site recalibrations are not proceeding correctly (or with the required frequency).
  • The EXO's turbidity sensor ceased to report after the 5-day outage in March, and may in fact have been removed.  The other sensors (blue-green algae, chlorophyll, fDOM) continue to fluctuate above and below zero and it is uncertain whether these data provide any value.
  • The analog AirT/RH sensor remains offline, as it has been since mid-2014. Based on my experience with a similar failure and repair at the Calabash Caye buoy I tend to believe that this is a problem with buoy wiring and not a failure of the RM Young sensor itself.
  • We do, however, have both air temperature and humidity reports from the integrated WXT unit. Although these reports cannot be validated by redundant readings from the analog AirT/RH sensor, they do appear to be believable.
  • Wind speeds and directions, as well as barometric pressures, continue to be reported redundantly by multiple sensors, and all appear to be consistent and correct.
  • The diagnostic RH (humidity) sensors within the buoy's two sealed junction boxes (JBs) show some cause for concern. The "Met" JB humidity levels remain acceptably low, although they have been climbing throughout the buoy's lifetime, and desiccant replacement might be helpful. Prior to the 12/8/2016 swapout the "Main" JB's humidity levels ranged between 60% and 90%, which is more or less the same as outside RHs and is unacceptably moist. After the swapout these RHs dropped to the 40% - 55% range, which is an improvement but still alarmingly high, and they have been increasing ever since.  Main JB RHs presently range from about 55% to 70%. The Main JB should be examined to ensure that it is free of cracks, that its o-ring is intact and properly seated, and that its desiccant is fresh.
We do not have any reports about how often the EXO Sonde's modules have been recalibrated on-site since December of 2015. According to the manufacturer many of these modules should be recalibrated on a monthly basis.

I will post any further updates that are received at AOML.  The next annual maintenance operation should be carried out this coming December, or earlier.

Mike J+

Friday, February 5, 2016

change in maintenance team for Tobago stations

[This blog update is back-dated to February 5, 2016, which is when AOML received the email announcement reproduced here. -- Mike Jankulak]
Good morning,

I write to informed you that effective March 3rd 2016,  Jahson Alemu would no longer be employed at the IMA.  Mr Jonathan Gomez and Mr Addison Titus (copied on this email) will be responsible for CREWS maintenance until we find a replacement for Mr Alemu.

If there are any concerns/ issues with regards to the CREWS in Tobago, can you please send an email to me and copy Mr Gomez and Mr Titus.

Thanks for your continued cooperation.

Best regards
Rahanna

Rahanna A. Juman, PhD
Principal Research Officer/ Deputy Director (Ag.)
Institute of Marine Affairs
Hilltop Lane, Chaguaramas
Trinidad and Tobago
West Indies

Monday, December 7, 2015

2015 Annual maintenance

Following servicing and repair by YSI (much appreciated), the sonde was received in Trinidad and Tobago in December 2015.

Over the period 7-8 December ARTO1 will undergo annual maintenance, which would involve bringing the unit to shore for a thorough inspection of all instruments and the re-installation of the sonde.

7th Dec 2015 (day 1)

No obvious damage was noted to any of the above water instruments. Fisher folk and other boaters continue not to moor on the unit.(....for now!!!). As expected the base and anchor ropes were laden with large colonies of barnacle, especially as the unit has not been visited since the EXO2 was removed for repair. The buoy base, ropes and sub surface floats did not seem to be compromised as a result of the fouling.

The unit was brought to shore, cleaned and inspected. Interface with the ARTO1 could only be established by serial connection. The RF and modem methods were unsuccessful. This seems to be a recurrent problem in interfacing with the units, especially for the purposes of data back-up, as logistically it is not always feasible to serial connect to the main data logger. (Ideas??)

The WXT was adjusted to align more with the RM Young anemometers, as recommended by Mike, This will be looked at again later tonight and before re-deployment tomorrow. The current disparity between the two units seems to be between 5-7 deg.

Some pooled water was noted within the main bulkhead. The O-ring was replaced for good measure, and desiccant is currently drying, to be put into the buoy tonight. RH levels will be checked in the morning.

No pictures for now, maybe tomorrow.


8th Dec 2015 (day 2)

We ran into some problems with the oven (not a drying oven, just a regular one), so I was not able to let the dessicant dry for as long as required prior to redeployment. I'll keep an eye on this as it is currently ~45% RH in the Main Buoy. If needed, in the near future I'll have to dry out the dessicant again and re-install into the buoy, however I will not bring the buoy in for this (it will be tricky).

The depressurization valve seems to be slowly corroding and may have to be changed at the next annual maintenance trip if not sooner. Again, I'll keep an eye on this.

Cell and RF connections were still unsuccessful, and serial connection was working fine. All data was downloaded and backed up.

All connections and ports were clean and unbroken.

The met data logger humidity was not checked due to an oversight. This will be monitored over the next few months.

The EXO2 was installed and all seemed functional. A new umbilical cord should be acquired to reduce the strain on the connector cable to the sonde. Also, some sacrificial anodes should be acquired. A dummy plug currently fills a vacant slot and the anode  on the outside of the buoy are near fully dissolved.

9th Dec 2015

Cellular connections were established after re-setting the Loggernet pathway.